# HeadGAP: Few-shot 3D Head Avatar via Generalizable Gaussian Priors

### v1

@@ 3.2 Avatar Representation
where $\mathcal{A}^{g}=\{\bm{\mu}^{\prime},\mathbf{r}^{\prime},\mathbf{S}^{\prime},\alpha\}$ is a point’s local attributes except from appearance items, and $\mathbf{e}\coloneqq\mathcal{D}(\bm{\mu}^{\prime})=\mathcal{T}(\bm{\mu}^{\prime})-\mathcal{T}(\bm{\mu}^{\prime}_{neutral})$ is the point-specific dynamic signals obtained by subtracting the global neutral point position $\mathcal{T}(\bm{\mu}^{\prime}_{neutral})$ from the global posed point position $\mathcal{T}(\bm{\mu}^{\prime})$. Both $f^{\mathcal{M}_{1}}_{p}$ and $f^{\mathcal{M}_{2}}_{p}$ are part-specific MLPs. We define the overall dynamic modeling as: $\mathcal{A}^{f}=f^{\mathcal{M}}_{p}(\mathbf{f},\mathbf{z})$ , where $\mathcal{A}^{f}=\mathcal{A}^{g}\cup\{\mathbf{h}\}$ denotes the final Gaussian attributes used for splatting.
@@
Gaussian Splatting with CNN refinement. Inspired by [74, 75], we apply a screen-space CNN $f^{\mathcal{C}}$ to refine the rendered results:
@@
Different from [74, 75], we do not conduct super-resolution, but keep the input and output with the same resolution for refinement. We aim to use large-scale training data to enable CNN to capture generalizable structured appearance priors that are challenging to exploit by our 3DGS-based representation. As indicated in Sec. 4.5, using CNN-based refinement can indeed capture those priors to render more photo-realistic results for few-shot personalization.
@@ 3.3 Head Prior Learning
where $\mathcal{L}_{rec}$ and $\mathcal{L}_{reg}$ denote the image reconstruction loss and training regularization loss respectively. The ground truth image is denoted as $\mathbf{I}^{*}$. To improve the fidelity of the mouth region, we further supervise masked rendered mouth region $\mathbf{I}_{m}$ with masked ground truth mouth region $\mathbf{I}^{*}_{m}$, inspired by [70]. Specifically, the image reconstruction loss:
@@ 3.4 Few-shot Personalization
Then, we start fine-tuning to update the network’s parameters so that the avatar can capture the details of the target identity from the inputs. We leverage prior knowledge in this procedure through three strategies. First, we use small learning rates for all parameters except $\mathbf{f}$. Next, we fully make use of extracted part-based priors by excluding the fine-tuning process for the mouth region, as modeling the highly flexible mouth region with few inputs is challenging. Finally, we apply view regularization to prevent overfitting to the target view, inspired by [83, 56]. Specifically, we constraint the fine-tuning results of some reference views with neutral face $\{\mathbf{R}_{i}\}_{i=1}^{m}$ to be close to the rendering results before fine-tuning $\{\tilde{\mathbf{R}}_{i}\}_{i=1}^{m}$, where $m$ is the number of the generated reference views. With the prior knowledge, our personalized avatar achieves stable reenactment while preserving the details of the target identity. The fine-tuning is conducted by minimizing the loss function in Sec. 3.3:
@@ 4.1 Setup
FLAME Tracking. Inspired by [54, 86], we designed a tracking algorithm that optimizes the FLAME  [43] parameters, which can adapt to different numbers of input views. In the following experimental section, unless otherwise stated, the training data uses this tracking algorithm to obtain the ground truth 3DMM parameters. For the test data, we simplify the tracking to use sparse view inputs and provide the corresponding 3DMM parameters for testing.
@@ A Implementation Details
Model Detail. We divide the Gaussian primitives into $p=11$ parts, including 1) “forehead”, 2) “nose”, 3) “eye”, 4) “teeth”, 5) “lip”, 6) “ear”, 7) “hair”, 8) “boundary”, 9) “neck”, 10) “other face region”, and 11) “other”. The part for the primitives is determined by the face masks provided by FLAME [43]. The illustration of Gaussian primitives with different parts is shown in Fig. A. The primitive number is set to $n=83,651$ by initializing from a UV map with a resolution of $300\times 300$. The feature dimensions of identity-shared point encoding $\mathbf{f}$, identity code $\mathbf{z}$, and point appearance feature $\mathbf{h}$ are set to $c_{1}=48$, $c_{2}=128$, and $c_{3}=34$ respectively. All the MLPs $f^{\mathcal{M}}$ consist of $4$ layers. Meanwhile, the CNN $f^{\mathcal{C}}$ contains $6$ layers.
Training Detail. We adopt Adam [39] optimizer for the model training. For prior learning, we utilize $k=119$ identities and set the batch size to $32$. For all the parameters, the learning rate begins at $1e^{-3}$ and decreases with the cosine scheduler. The prior model is trained on $8$ A100 GPUs for 100K steps, which takes around 2 days. The loss weights $\lambda_{m}$, $\lambda_{l1}$, $\lambda_{ssim}$, $\lambda_{lpips}$, $\lambda_{\alpha}$, $\lambda_{s}$, $\lambda_{\mu}$, and $\lambda_{arap}$ are set to $10$, $0.8$, $0.2$, $0.4$, $1$, $1$, $0.01$, and $1$ respectively. For few-shot personalization, we set the batch size to $1$. We set the learning rate of the identity-shared point encoding $\mathbf{f}$ to $1e^{-3}$ and other parameters’ to $1e^{-5}$. Unless otherwise stated, we take $500$ steps for inversion and $500$ steps for fine-tuning, which uses about $5$ minutes in total with an A100 GPU. For view regularization, we generate $m=16$ reference views similar to the camera setups of the NeRSemble dataset. The loss weights $\lambda_{ref}$ is set to $0.01$. For 3-shot novel identities’ personalization on NeRSemble, we utilize cameras with id “0”, “8”, and “15”. For our captured data, we select viewpoints similar to NeRSemble.
@@ B.2 Prior Learning Results
We show our prior learning results of the 119 identities in Fig. G. The visualized results show that our GAPNet can learn the appearance characteristics of different identities.
@@ B.3 More Qualitative Results
Fig. C includes more self-reenactment visualization results for different identities. We showcase the novel-view rendering results of the personalized head avatars. Fig. D includes more cross-reenactment results for different identities from the frontal view, which showcases the stable animations of our approach. Fig. E includes additional cross-reenactment visualization results for different identities for some side views.
@@ C.3 Comparison with GS-based methods
We show the comparison with GaussianAvatars [54] and FlashAvatar [70] in Fig. I. Although they do not focus on few-shot input like ours, we include comparisons because we all use 3D Gaussian Splatting as a representation. We observe that they require a substantial overlap of input views or monocular videos with human heads rotated to different orientations.
In few-shot personalization experiments, as shown in the Fig. I, all per-subject optimization Gaussian Splatting-based baselines lack prior information and require individual training for each person. It can be observed that all baseline methods tend to overfit the training views and fail to extrapolate to unseen views. This qualitative comparison demonstrates the effectiveness and necessity of constructing priors for Gaussian Splatting. Due to the noticeable artifacts, FlashAvatar, GaussianAvatars, and GaussianAvatars are infeasible for calculating meaningful ID similarity metrics. Therefore, we did not report their corresponding metrics in Tab.1 of the main text.
@@ D Limitations and Future Works
While our method can quickly construct personalized, high-fidelity, and realistic human head avatars, it still has the following issues: 1. In cases where the subject wears glasses or has noticeable facial accessories, the avatar construction may exhibit artifacts (as depicted in Fig. J). A reason for this incapability is that our prior learning phase does not incorporate such samples for training. Including the corresponding data for training can potentially resolve this problem. 2. Our method does not focus on modeling the subject’s clothing and hair. We believe that combining methods such as [47] to model hair or clothing separately is a promising research direction. 3. Additionally, lighting variation is important for the realism of head avatars. Currently, we only consider uniform lighting. We believe that integrating relighting into the Gaussian Splatting is also a promising research direction for head avatars.
@@
Figure G: The rendered results of the 119 identities used for prior learning.
Figure H: More qualitative experiments on other subjects using 3-shot inputs compared to state-of-the-art methods.
Figure I: Qualitative comparison results. We compare the rendering results from different views using our 3-shot input avatars with the Gaussian Splatting-based baseline methods.
Figure J: Failure cases. Our approach can not resolve subjects with noticeable facial accessories (e.g., glasses).

### v2

@@ 3.2 Avatar Representation
where $\mathbf{h}$ is point appearance attribute, $\mathcal{A}^{g}=\{\bm{\mu}^{\prime},\mathbf{r}^{\prime},\mathbf{S}^{\prime},\alpha\}$ denotes other attributes, and $\mathbf{e}\coloneqq\mathcal{D}(\bm{\mu}^{\prime})=\mathcal{T}(\bm{\mu}^{\prime})-\mathcal{T}(\bm{\mu}^{\prime}_{neutral})$ is the point-specific dynamic signals obtained by subtracting the global neutral point position $\mathcal{T}(\bm{\mu}^{\prime}_{neutral})$ from the global posed point position $\mathcal{T}(\bm{\mu}^{\prime})$. Both $f^{\mathcal{M}_{1}}_{p}$ and $f^{\mathcal{M}_{2}}_{p}$ are part-specific MLPs. We define the overall dynamic modeling as: $\mathcal{A}^{f}=f^{\mathcal{M}}_{p}(\mathbf{f},\mathbf{z})$ , where $\mathcal{A}^{f}=\mathcal{A}^{g}\cup\{\mathbf{h}\}$ denotes the final Gaussian attributes used for splatting.
@@
Gaussian Splatting with CNN refinement. Inspired by recent works [74, 75], we apply a screen-space CNN $f^{\mathcal{C}}$ to refine the rendered results:
@@
Different from previous methods [74, 75], we do not conduct super-resolution, but keep the input and output with the same resolution for refinement. We aim to use large-scale data to enable CNN to capture generalizable structured appearance priors that are challenging to exploit by our 3DGS-based representation. As indicated in Sec. 4.5, using CNN-based refinement can indeed capture those priors to render more photo-realistic results for few-shot personalization.
@@ 3.3 Head Prior Learning
where $\mathcal{L}_{rec}$ and $\mathcal{L}_{reg}$ denote the image reconstruction loss and training regularization loss respectively. The ground truth image is denoted as $\mathbf{I}^{*}$. To improve the fidelity of the mouth region, we further supervise mouth region $\mathbf{I}_{m}$ with masked ground truth mouth region $\mathbf{I}^{*}_{m}$, inspired by FlashAvatar [70]. Specifically, the image reconstruction loss:
@@ 3.4 Few-shot Personalization
Then, we start fine-tuning to update the network’s parameters so that the avatar can capture the details of the target identity from the inputs. We leverage prior knowledge in this procedure through three strategies. First, we use small learning rates for all parameters except $\mathbf{f}$. Next, we exploit extracted part-based priors by excluding the fine-tuning for the mouth region, as modeling the highly flexible mouth region with few inputs is challenging. Finally, we apply view regularization to prevent overfitting to the target view, inspired by previous methods [83, 56]. Specifically, we constraint the fine-tuning results of some reference views with neutral face $\{\mathbf{R}_{i}\}_{i=1}^{m}$ to be close to the rendering results before fine-tuning $\{\tilde{\mathbf{R}}_{i}\}_{i=1}^{m}$, where $m$ is the number of the generated reference views. With the prior knowledge, our personalized avatar achieves stable reenactment while preserving the details of the target identity. The fine-tuning is conducted by minimizing the loss function in Sec. 3.3:
@@ 4.1 Setup
FLAME Tracking. Inspired by previous studies [54, 86], we designed a tracking algorithm that can optimize the FLAME  [43] parameters with different numbers of input views. In the following experimental section, unless otherwise stated, the training data uses this tracking algorithm to obtain the ground truth 3DMM parameters. For the test data, we simplify the tracking to use sparse view inputs and provide the corresponding 3DMM parameters for testing.
@@ A Implementation Details
Model Detail. We divide the Gaussian primitives into $p=11$ parts, including 1) “forehead”, 2) “nose”, 3) “eye”, 4) “teeth”, 5) “lip”, 6) “ear”, 7) “hair”, 8) “boundary”, 9) “neck”, 10) “other face region”, and 11) “other”. The part for the primitives is determined by the face masks provided by FLAME [43]. The illustration of Gaussian primitives with different parts is shown in Fig. A. The primitive number is set to $n=83,651$ by initializing from a UV map with a resolution of $300\times 300$. The feature dimensions of identity-shared point encoding $\mathbf{f}$, identity code $\mathbf{z}$, and point appearance feature $\mathbf{h}$ are set to $c_{1}=48$, $c_{2}=128$, and $c_{3}=34$ respectively. All the MLPs $f^{\mathcal{M}}$ consist of $4$ layers. Meanwhile, the CNN $f^{\mathcal{C}}$ contains $6$ layers. The identity codebook $\mathbf{z}$ is initialized with zero.
Training Detail. We adopt Adam [39] optimizer for the model training. For prior learning, we utilize $k=119$ identities and set the batch size to $32$. For all the parameters, the learning rate begins at $1e^{-3}$ and decreases with the cosine scheduler. The prior model is trained on $8$ A100 GPUs for 100K steps, which takes around 2 days. The loss weights $\lambda_{m}$, $\lambda_{l1}$, $\lambda_{ssim}$, $\lambda_{lpips}$, $\lambda_{\alpha}$, $\lambda_{s}$, $\lambda_{\mu}$, and $\lambda_{arap}$ are set to $10$, $0.8$, $0.2$, $0.4$, $1$, $1$, $0.01$, and $1$ respectively. For few-shot personalization, we set the batch size to $1$. We set the learning rate of the identity-shared point encoding $\mathbf{f}$ to $1e^{-3}$ and other parameters’ to $1e^{-5}$. Unless otherwise stated, we take $500$ steps for inversion and $500$ steps for fine-tuning, which uses about $5$ minutes in total with an A100 GPU. For view regularization, we generate $m=16$ reference views similar to the camera setups of the NeRSemble dataset. The loss weights $\lambda_{ref}$ is set to $0.01$. For 3-shot novel identities’ personalization on NeRSemble, we utilize cameras with id “0”, “8”, and “15”. For our captured data, we select viewpoints similar to those of NeRSemble.
@@ B.2 Prior Learning Results
We show our prior learning results of the 119 identities in Fig. H. The visualized results show that our GAPNet can learn the appearance characteristics of different identities.
@@ B.3 More Qualitative Results
Fig. C shows self-reenactment results with novel-view renderings for different identities. Fig. D and Fig. E present cross-reenactment results from frontal and side views, respectively, demonstrating stable animations.
The results indicate that our model effectively generalizes to data that differs from the NeRSemble dataset. Furthermore, it achieves consistent few-shot performance across diverse ethnicities and genders, thereby further reinforcing its capacity for effective generalization.
@@ B.4 Head Avatar Editing
B.5 More In-the-wild Results
In this section, we present additional results on in-the-wild images. All result IDs are out-of-distribution samples beyond the NeRSemble [40] dataset. Specifically, we capture monocular video data of each identity performing various expressions and select 12 images for avatar personalization. As shown in  Fig. G, we present the cross-reenactment driving results when providing the same facial expression motion sequence. The results demonstrate that our method exhibits strong few-shot generalization capability even in in-the-wild settings.
@@ C.3 Comparison with GS-based methods
We show the comparison with GaussianAvatars [54] and FlashAvatar [70] in Fig. J. Although they do not focus on few-shot input like ours, we include comparisons because we all use 3D Gaussian Splatting as a representation. We observe that they require a substantial overlap of input views or monocular videos with human heads rotated to different orientations.
In few-shot personalization experiments, as shown in the Fig. J, all per-subject optimization Gaussian Splatting-based baselines lack prior information and require individual training for each person. It can be observed that all baseline methods tend to overfit the training views and fail to extrapolate to unseen views. This qualitative comparison demonstrates the effectiveness and necessity of constructing priors for Gaussian Splatting. Due to the noticeable artifacts, FlashAvatar, GaussianAvatars, and GaussianAvatars are infeasible for calculating meaningful ID similarity metrics. Therefore, we did not report their corresponding metrics in Tab.1 of the main text.
@@ D Limitations and Future Works
While our method can quickly construct personalized, high-fidelity, and realistic human head avatars, it still has the following issues: (1) In cases where the subject wears glasses or has noticeable facial accessories, the avatar construction may exhibit artifacts (as depicted in Fig. K). A reason for this incapability is that our prior learning phase does not incorporate such samples for training. Including the corresponding data for training can potentially resolve this problem. (2) The adoption of CNNs for refinement in screen space may result in view-dependent overfitting, which can induce flickering among different viewpoints and lead to quality degradation for certain unseen views during training. Therefore, exploring more consistent refinement techniques in 3D space presents a promising avenue for further investigation. (3) Our method does not focus on modeling the subject’s clothing and hair. We believe that combining methods such as [47] to model hair or clothing separately is a promising research direction. (4) Additionally, lighting variation is important for the realism of head avatars. Currently, we only consider uniform lighting. We believe that integrating relighting into the Gaussian Splatting is also a promising research direction for head avatars.
@@
Figure G: Qualitative results of 3D animatable head avatars generated from few-shot in-the-wild images and driven by the same facial expression sequence.
Figure H: The rendered results of the 119 identities used for prior learning.
Figure I: More qualitative experiments on other subjects using 3-shot inputs compared to state-of-the-art methods.
Figure J: Qualitative comparison results. We compare the rendering results from different views using our 3-shot input avatars with the Gaussian Splatting-based baseline methods.
Figure K: Failure cases. Our approach can not resolve subjects with noticeable facial accessories (e.g., glasses).